# Principles for Optimizing Quantum Transduction in Piezo-Optomechanical Systems

### v2

@@ VI Conclusion
With these fundamental ideas in mind, we examined how the transduction efficiency changes as a function of these interaction constants using the nominal system parameters in [16] as a starting point. We found that increasing both the electromechanical coupling and waveguide coupling will increase the maximum transduction efficiency, albeit requiring a substantially higher power for the driving laser. However, by increasing the optomechanical coupling as well, we can decrease the required power of the driving laser to more reasonable values to achieve the maximum efficiency. In achieving these efficiencies, we also examined over what bandwidth we may transduce light at high efficiency, so that we may understand with what fidelity single-photon pulses of finite bandwidth can be transduced. Defining the transduction bandwidth as the range over which the transduction efficiency is greater than $50\%$ of its maximum, we find that this bandwidth can be enhanced substantially by increasing the coupling strength of the electromechanical and waveguide interactions. This comes with a frequency shift that must be accounted for when selecting the correct frequency of the driving laser. Among other things, we also found that it is critical to use a driving laser with linewidth much narrower than the GHz-scale microwave frequencies being transduced in order to avoid two-mode squeezing generating light at the transduction frequency that is not part of the transduction process. With an optimum choice of materials, systems of this type will work well as transducers, provided the light they generate can be well-separated from the driving laser field. While this frequency separation problem at the single-photon level is difficult to address using ordinary optics for an ordinary optomechanical coupling, we also show how this challenge may be solved by instead using a new optomechanical interaction utilizing higher orders of photoelasticity and electrostriction.
@@ Acknowledgements.
We gratefully acknowledge support from our colleagues at the Air Force Research Laboratory, at Purdue University, and in particular, the insightful comments and discussions with Drs. Michael Senatore and Dylan Heberle. The views expressed are those of the authors and do not reflect the official guidance or position of the United States Government, the Department of Defense or of the United States Air Force. The appearance of external hyperlinks does not constitute endorsement by the United States Department of Defense (DoD) of the linked websites, or of the information, products, or services contained therein. The DoD does not exercise any editorial, security, or other control over the information you may find at these locations.
@@ Appendix A Table of material parameters
Material $e_{31}$ ($C/m^{2}$) $\epsilon_{33}$ $\rho$ (in $g/cm^{3}$) $p_{3333}$ fab. [EM]$\;\;\frac{e_{31}}{\rho\epsilon_{33}}$ [OM]$\;\;\frac{\epsilon_{33}p_{3333}}{\sqrt{\rho}}$
$\text{Al}_{2}\text{O}_{3}$ (sapphire) 0 (c.s.) [48] 8.6 [49] 3.97,3.99 $(\gamma,\alpha)$[50] $-0.2$ [50] FC 0 $\approx-0.86\times 10^{0}$
$\text{Al}_{2}\text{O}_{3}$ (alumina) 0 (c.s.) 11.54 [50] 3.8 [50] $-0.237$[51] (calc) $\checkmark$ 0 $\approx-1.40\times 10^{0}$
AlN 0.58 [48] 10.7 [52] 3.255 [50] -0.107 [53] $\checkmark$ $1.67\times 10^{-2}$ $\approx-0.63\times 10^{0}$
$\text{BaB}_{2}\text{O}_{4}$ (BBO) 0.092 [54] (calc) $\approx 7.65$ [55] 3.85 [56] 0.039 [57] X $3.12\times 10^{-3}$ $\approx 0.15\times 10^{0}$
$\text{BaTiO}_{3}$ 2.16 [48] 2000 [49] 5.9988-6.02 [50] 0.77 [58] $\checkmark$ $\approx 1.8\times 10^{-4}$ $\approx 6.28\times 10^{2}$
$\text{In}_{2}\text{O}_{3}+\text{Sn}\text{O}_{2}$ (ITO) 0 (c.s.) [59, 60] $\approx 4$ [61] 6.8 [62] ? $\checkmark$ 0 ?
$\text{KD}_{2}\text{PO}_{4}$ (DKDP) 0 (pz) [63, 64] 50 [63] 2.344 [63] 0.17 [65] X 0 $\approx 5.55\times 10^{0}$
$\text{KH}_{2}\text{PO}_{4}$ (KDP) 0 (pz) [66, 64] 44.4 [49] 2.338 [56] 0.122 [50] X 0 $\approx 3.54\times 10^{0}$
$\text{KTiOPO}_{4}$ (KTP) 0.5428 [67] $>17.5$ [68] 3.024 [56] ? FC $>1.026\times 10^{-2}$
$\text{KTa}_{1-x}\text{Nb}_{x}\text{O}_{3}$ (KTN) -5.2 [69] $\approx 435$ [69] 5.759 [69] ? FC $-2.08\times 10^{-3}$
$\text{LiB}_{3}\text{O}_{5}$ (LBO) 0.63 [70] 10.71 [71] 2.474 [56] 0.3 [72] FC $\approx 2.38\times 10^{-2}$ $\approx 2.04\times 10^{0}$
$\text{LiNbO}_{3}$ $0.37$ [48] $30$ [50] 4.3-4.7 [50] $0.118$ [73] FC $\approx 2.74\times 10^{-3}$$\approx\pm 1.67\times 10^{0}$
$\text{LiTaO}_{3}$ $\approx 0$ [74] $46$ [50] 7.45 [50] -0.044 [50] FC $\approx 0$ $\approx-0.74\times 10^{0}$
$\text{PbZr}_{1-x}\text{Ti}_{x}\text{O}_{3}$ (PZT) $\approx-4.53$ (mean) [75] $\{1440,3030\}$ [76] 7.5 [77] ???? FC $\{1.9,4.2\}\times 10^{-4}$???
Si (crystal) 00 (c.s.) $12.1$* [50] 2.3296 [50] -0.097 (calc) [78] $\checkmark$ 0 $\approx 0.77\times 10^{0}$
Si (amorphous) 0 $\approx 11.9$ [79, 50](calc) 2.288 [79] ? $\checkmark$ 0 ?
$3C-\text{SiC}$ 0 (pz) [80] 9.72 [81] 3.21 [81] -0.110 [78](calc) $\checkmark$ 0 $\approx-0.59\times 10^{0}$
$4H-\text{SiC}$0.0623 [80] 10.32 [81] 3.21 [81] ? $\checkmark$ $-1.88\times 10^{-3}$
$6H-\text{SiC}$0.1158 [80] 10.03 [81] 3.21 [81] ? $\checkmark$ $-3.6\times 10^{-3}$
$\text{Si}_{3}\text{N}_{4}$ 00 (c.s.) $4.2$ [50] 3.17 [50] 0.239 [16] $\checkmark$ 0 $0.564\times 10^{0}$
$\text{SiO}_{2}$ (amorphous) 0 (c.s.) 3.75[50] 2.196 [50] $0.121$[50] $\checkmark$ 0 $0.31\times 10^{0}$
$\alpha-\text{SiO}_{2}$ (quartz) 0 [82, 83, 26] (pz) $4.60$ [50] 2.648 [50] 0.1 [50] X 0 $\approx 0.28\times 10^{0}$
@@
$\beta-\text{Ta}_{2}\text{O}_{5}$ 0 [85] (pz) 19.5 [86] 8.015 [87] $\approx-0.05$ (est)[88] $\checkmark$ 0 $\approx-0.34\times 10^{0}$
ZnO $0.53$ [48] $8.2$ [50] 5.6-5.676 [50] $-0.235$ [50] X $\approx 1.15\times 10^{-2}$ $\approx-0.81\times 10^{0}$

### v4

@@ V Conclusion
With these fundamental ideas in mind, we examined how the transduction efficiency changes as a function of these interaction constants using the nominal system parameters in [16] as a starting point. We found that increasing both the electromechanical coupling and waveguide coupling will increase the maximum transduction efficiency, albeit requiring a substantially higher power for the driving laser. However, by increasing the optomechanical coupling as well, we can decrease the required power of the driving laser to more reasonable values to achieve the maximum efficiency. In achieving these efficiencies, we also examined over what bandwidth we may transduce light at high efficiency, so that we may understand with what fidelity single-photon pulses of finite bandwidth can be transduced. Defining the transduction bandwidth as the range over which the transduction efficiency is greater than $50\%$ of its maximum, we find that this bandwidth can be enhanced substantially by increasing the coupling strength of the electromechanical and waveguide interactions. This comes with a frequency shift that must be accounted for when selecting the correct frequency of the driving laser. Among other things, we also found that it is critical to use a driving laser with linewidth much narrower than the GHz-scale microwave frequencies being transduced in order to avoid two-mode squeezing generating light at the transduction frequency that is not part of the transduction process. With an optimum choice of materials, systems of this type will work well as transducers, provided the light they generate can be well-separated from the driving laser field.
@@ Acknowledgements.
We gratefully acknowledge support from our colleagues at the Air Force Research Laboratory, at Purdue University, and in particular, the insightful comments and discussions with Drs. Michael Senatore and Dylan Heberle, as well as an anonymous referee. The views expressed are those of the authors and do not reflect the official guidance or position of the United States Government, the Department of Defense or of the United States Air Force. The appearance of external hyperlinks does not constitute endorsement by the United States Department of Defense (DoD) of the linked websites, or of the information, products, or services contained therein. The DoD does not exercise any editorial, security, or other control over the information you may find at these locations.
@@ Appendix A Table of material parameters
Material $h_{33}$ ($V/m$) $\epsilon_{33}$ (RF) $\epsilon_{33}$ (IR) $\rho$ ($g/cm^{3}$) $p_{3333}\rightarrow p_{33}$ fab. [EM]$h_{33}\!\sqrt{\!\frac{\epsilon_{33}(\!\text{RF}\!)}{\rho}}$ [OM]$\frac{\!\epsilon_{33}\!(\!\text{IR}\!)p_{33}}{\sqrt{\rho}}$
$\text{Al}_{2}\text{O}_{3}$ (sapphire) 0 (c.s.) [50] 8.6 [51] 3.05 [52] 3.99 [53] $-0.2$ [53] FC 0 $-3.06\times 10^{-1}$
$\text{Al}_{2}\text{O}_{3}$ (alumina) 0 (c.s.) 11.54 [53] 2.72 [54] 3.8 [53] $-0.237$[55] (clc) $\checkmark$ 0 $-3.31\times 10^{-1}$
AlN 0.145 [50] 9.5 [56] 3.67 [57] 3.255 [53] -0.107 [58] $\checkmark$ $2.47\times 10^{-1}$ $-2.18\times 10^{-1}$
$\text{BaB}_{2}\text{O}_{4}$ (BBO) 0.0224 [59] (clc) $\approx 7.65$ [60] 2.37 [61] 3.85 [62] 0.039 [63] X $3.16\times 10^{-2}$ $5.39\times 10^{-2}$
$\text{BaTiO}_{3}$ 0.0017 [50] 2000 [51] 4.88 [64] $\approx 6.00$(m)[53] 0.77 [65] $\checkmark$ $3.15\times 10^{-2}$ $1.53\times 10^{0}$
$\text{\text{In}}_{2}\text{O}_{3}\!+\!\text{Sn}\text{O}_{2}$(ITO) 0 (c.s.) [66, 67] $\approx 4$ [68] $\approx 3.6$ [69] 6.8 [70] ? $\checkmark$ 0 ?
$\text{KD}_{2}\text{PO}_{4}$ (DKDP) 0 (pz) [71, 72] 50 [71] (op)[62] 2.344 [71] 0.17 [73] X 0 (op)
$\text{KH}_{2}\text{PO}_{4}$ (KDP) 0 (pz) [74, 72] 44.4 [51] (op) 2.338 [62] 0.122 [53] X 0 (op)
$\text{KTiOPO}_{4}$ (KTP) 0.0935 [75] $>17.5$ [76] 3.297 [77] 3.024 [62] ? FC $2.25\times 10^{-1}$
$\text{KTa}_{1\!-\!x}\text{Nb}_{x}\!\text{O}_{3}$(KTN) 0.0267 [78] $\approx 435$ [78] $\approx 4.74^{**}$[79] 5.759 [78] ? FC $2.32\times 10^{-1}$
$\text{LiB}_{3}\text{O}_{5}$ (LBO) 0.0971 [80] 10.71 [81] $\approx 2.5^{***}$[82] 2.474 [62] 0.3 [83] FC $2.02\times 10^{-1}$ $4.77\times 10^{-1}$
$\text{LiNbO}_{3}$ 0.0433 [50] $30$ [53] 4.569 [84] 4.3-4.7 [53] $0.118$ [85] FC $1.12\times 10^{-1}$ $2.54\times 10^{-1}$
$\text{LiTaO}_{3}$ 0.0413 [86] $46$ [53] 4.506 [87] 7.45 [53] -0.044 [53] FC $1.03\times 10^{-1}$ $-7.26\times 10^{-2}$
$\text{PbZr}_{1-x}\!\text{Ti}_{x}\text{O}_{3}$(PZT) $8.68\!\times\!10^{-3}\!$(m)[88] $\approx\!2200$(m)[89] 5.76 [90] 7.5 [91] ???? FC $1.49\times 10^{-1}$???
Si (crystal) 00 (c.s.) $12.1$* [53] 12.1 [62] 2.330 [53] -0.097 (clc) [92] $\checkmark$ 0 $-7.69\times 10^{-1}$
Si (amorphous) 0 $\approx 11.9$ [53](clc) 12.1 [93] 2.288 [94] ? $\checkmark$ 0 ?
$3C-\text{SiC}$ 0 (pz) [95] 9.72 [96] 6.6 [97] 3.21 [96] -0.11 [92](clc) $\checkmark$ 0 $-4.05\times 10^{-1}$
$4H-\text{SiC}$ $-6.04\!\times\!10^{-3}$ [95] 10.32 [96] 6.81 [98] 3.21 [96] ? $\checkmark$ $-1.08\times 10^{-2}$
$6H-\text{SiC}$ $-1.16\!\times\!10^{-2}$ [95] 10.03 [96] 6.57 [98] 3.21 [96] ? $\checkmark$ $-2.04\times 10^{-2}$
$\text{Si}_{3}\text{N}_{4}$(amorphous) 00 (c.s.) $4.2$ [53] 3.99 [99] $\approx 3.17$ [53] 0.239 [16] $\checkmark$ 0 $5.35\times 10^{-1}$
$\text{SiO}_{2}$ (amorphous) 0 (c.s.) 3.75[53] 2.09 [100] 2.196 [53] $0.121$[53] $\checkmark$ 0 $1.45\times 10^{-1}$
$\alpha-\text{SiO}_{2}$ (quartz) 0 [101, 102, 24] (pz) $4.60$ [53] 2.36 [103] 2.648 [53] 0.1 [53] X 0 $1.45\times 10^{-1}$
@@
$\beta-\text{Ta}_{2}\text{O}_{5}$ 19.5 [106] 4.20 [107] 8.015 [108] $\approx\!-0.05\!$ (est)[109] $\checkmark$ ? $-7.41\times 10^{-2}$
ZnO $0.0906$ [50] $8.2$ [53] 3.01 [110] 5.6-5.676 [53] $-0.235$ [53] X $1.09\times 10^{-1}$ $-2.98\times 10^{-1}$